# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v1

@@ 5.3. Families of immersed atomic Lagrangians
Let $p\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to C$ and $q\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to C$ be the relative moduli spaces, which are proper. We denote by $r_{c}\colon M^{\mathcal{Y}_{c}}_{\underline{\sigma^{\prime}}_{c}}(a,b)\to M^{\mathcal{X}_{c}}_{\underline{\sigma}_{c}}(a,b)$ the morphism in Theorem 3.7 for each $c\in C$. By assumption, the push-forward of $\mathcal{Y}\to\mathcal{X}$ and the relative projection functor $\mathrm{D}^{\mathrm{b}}(\mathcal{X})\to\mathcal{K}u(\mathcal{X})$ induce a morphism $r^{\prime}\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ such that $r^{\prime}_{c}=r_{c}$. Therefore, $r^{\prime}$ is proper and quasi-finite, hence finite. Moreover, by [24, Proposition 5.1(1)], the fibers of $p$ are equidimensional of the same dimension and $M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)$ is also equidimensional. By Proposition 4.4, this implies that $r_{0*}\operatorname{\mathcal{O}}_{M^{Y}_{\sigma_{Y}}(a,b)}$ is atomic because $r_{1*}\operatorname{\mathcal{O}}_{M_{\underline{\sigma}_{1}}^{Y^{\prime}}(a,b)}$ is atomic. ∎
@@ 6.1. Double dual EPW sextics
Recall that by [16, Theorem 3.10], each smooth GM variety $X$ of dimension $\dim X\geq 3$ is uniquely determined by its Lagrangian data
@@
as in [50, Section 2]. See also [16, Appendix B] for an overview. We set $\mathsf{Y}_{A(X)}:=\mathsf{Y}_{A(X)}^{\geq 1}$ and $\mathsf{Y}_{A(X)^{\perp}}:=\mathsf{Y}_{A(X)^{\perp}}^{\geq 1}$, which are called the Eisenbud–Popescu–Walter (EPW) sextic and the dual EPW sextic associated with $X$, respectively.
Moreover, by [52, Section 1.2] and [18, Theorem 5.2], as $A(X)$ does not contain decomposable vectors [^6] This means $\mathbb{P}(A(X))\cap\operatorname{Gr}(3,V_{6})=\varnothing$, which holds for $A(X)$ when $X$ is smooth and $\dim X\geq 3$ by [16, Theorem 3.16]., we have natural double coverings
@@
Since $X$ and $Y$ are both ordinary, by [16, Propoition 3.14], the Lagrangian data of $X$ and $Y$ are $(V_{6},V_{5},A(X))$ and $(V_{6},V_{5},A(Y))$ respectively, with $\dim(A(X)\cap A(Y))=9$. Therefore, part (1) follows directly from the above definition of EPW varieties.
@@
Consequently, $\iota$ is finite and unramified in codimension one. Thus, the natural map $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is also unramified in codimension one. Then, by the smoothness of $\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ and [28, Theorem 1.6], it follows that $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified, and so is $\iota$.
Moreover, since $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified and finite of degree two, by [64, Tag 0B3D] and [64, Tag 00R4], it is flat and hence an étale double covering. In particular, $W$ is smooth. Then, by the uniqueness part of [18, Theorem 5.2], we have $W\cong\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and (2) follows. ∎
@@
It is suggested in [34, Section 5.1] that $i$ is injective, implying that $i$ is an embedding. Nevertheless, this is equivalent to $\varnothing=\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}\cap\mathsf{Y}_{A(X)^{\perp}}^{\geq 2}\subset\mathsf{Y}_{A(X)^{\perp}}$, which is impossible as explained in [23, Lemma 3.2]. See Remark 6.15 for an alternative proof of the non-injectivity of $i$.
@@ 6.1.1. GM threefolds in a fixed GM fourfold
constructed in [46] and [14]. As explained in [34, Section 5.1] (see also [21]), we have a commutative diagram
@@
where the left column is the natural inclusion and the right column is the finite and unramified morphism in Lemma 6.2(2). Using this relation, the following result is obtained in [34, Proposition 5.2].
@@
Since $X$ is very general, by [24, Corollary B.9], there is a commutative diagram
@@
As $X$ and $Y$ are general, we know that $F(X)$ and $F(Y)$ are both smooth by [34, Theorem 3.2] and [46]. Moreover, from [16, Theorem B.2], we may assume that
@@
then $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ are smooth by [18, Theorem 5.2]. Consequently, according to Proposition 6.4, $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an immersed Lagrangian submanifold.
Recall that the moduli stacks of ordinary GM varieties of dimension $n=3$ or $4$ are smooth irreducible separated Deligne–Mumford stacks of finite type over $\mathbb{C}$ (cf. [20, Corollary 5.12] and [38, Proposition A.2]). Since the locus of Hodge-special GM fourfolds is the union of countably many divisors by [15, Corollary 4.6], using Lemma 6.9, we can construct a smooth family of ordinary GM fourfolds $\mathcal{X}\to C$ over a smooth connected curve $C$ and a smooth divisor $\mathcal{Y}\subset\mathcal{X}$ such that the fibers over a point $0\in C$ are $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$. Moreover, there exists an uncountable set $F\subset C$ such that $\mathcal{X}_{c}$ is very general in the sense of Lemma 6.6 for each $c\in F$. Indeed, we start by considering a curve $C^{\prime}$ along with a family $\mathcal{X}^{\prime}\to C^{\prime}$ that satisfies the above assumptions. Next, we lift $C^{\prime}$ to a curve $C$ in the relative Hilbert scheme of hyperplane sections of $\mathcal{X}^{\prime}\to C^{\prime}$, obtaining the family $\mathcal{Y}\subset\mathcal{X}\to C$.
@@
Applying [20, Theorem 3.7, Proposition 4.1] to $\mathcal{X}\to C$ and $\mathcal{Y}\to C$, we obtain two families of Lagrangian data $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{X}))$ and $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{Y}))$ over $C$ [^7] See [20, Definition 3.15] for the precise definition., where $\mathscr{V}_{6}$, $\mathscr{V}_{5}$, $\mathscr{A}(\mathcal{X})$, and $\mathscr{A}(\mathcal{Y})$ are all locally free sheaves on $C$ such that
@@
for each $c\in C$. After replacing $C$ with an open dense neighborhood of $0\in C$, we can assume that these bundles are all trivial. Therefore, as in the proofs of [18, Theorem 5.2] and [17, Proposition 5.27], we get a family $p\colon\widetilde{\mathscr{Y}}_{\mathcal{X}}\to C$ of double dual EPW sextics and a family $q\colon\widetilde{\mathscr{Y}}^{\geq 2}_{\mathcal{Y}}\to C$ of double dual EPW surfaces such that $p^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ and $q^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}$ for each $c\in C$ by replacing $C$ with an open dense neighborhood of $0\in C$. Note that by our assumptions of families $\mathcal{X}$ and $\mathcal{Y}$, both $p$ and $q$ are smooth. We denote by $\mathscr{Y}_{\mathcal{X}}\to C$, $\mathscr{Y}_{\mathcal{Y}}^{\geq 2}\to C$, and $\mathscr{Y}^{\geq 2}_{\mathcal{X}}\to C$ the corresponding families of dual EPW sextics and dual EPW surfaces, respectively. Then $\mathscr{Y}_{\mathcal{X}},\mathscr{Y}^{\geq 2}_{\mathcal{X}}$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}$ are smooth. Note that they are all closed subschemes of $\mathbb{P}(\mathscr{V}^{\vee}_{6})\cong\mathbb{P}(V^{\vee}_{6})\times C$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}_{c}}\subset\mathscr{Y}_{\mathcal{X}_{c}}$ for each $c\in C$, we see $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}\subset\mathscr{Y}_{\mathcal{X}}$. Therefore, by passing to double coverings, taking simultaneous normalization (cf. [37, Theorem 12]), and using a relative version of Lemma 6.2, we have a commutative diagram
@@
Up to now, we have shrunk $C$ near $0\in C$ finitely many times. As $F$ is an infinite set, there is a point $t\in C\cap F$ such that Lemma 6.6 can be applied to $\mathcal{Y}_{t}\hookrightarrow\mathcal{X}_{t}$, i.e. $j_{t}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{t})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{t})^{\perp}}$ is an immersed atomic Lagrangian submanifold. Therefore, using Proposition 4.4, we know that the Lagrangian submanifold $i_{c}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ is atomic for each $c\in C$, and in particular, $j_{0}=i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is atomic. Finally, the computation of the cohomology group is done in [19, Proposition 2.5] and [46, Proposition 0.5], and the orthogonal complement of the kernel of the restriction map is computed in Corollary 5.2 and Remark 3.10. ∎
@@
To this end, we first assume that $X$ is projective. Let $l\colon X\hookrightarrow\mathbb{P}^{n}$ be an embedding, then each hyperplane $H\subset\mathbb{P}^{n}$ intersects with each divisor in $\{D_{i}\}_{i\in\mathbb{N}}$. By [42, Theorem 3.3.1], for each $t\in\mathbb{N}$, there exists an open dense subset $U_{t}\subset|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ such that when $[H]\in U_{t}$, $H\cap X$ and $H\cap D_{i}$ are irreducible and $H$ intersects with $D_{i}$ and $D_{i}\cap D_{j}$ properly for each $i,j\leq t,i\neq j$. Thus, $\dim H\cap D_{i}\cap D_{j}\leq\dim X-3$ for $i,j\leq t,i\neq j$. We define $U:=\cap_{t\in\mathbb{N}}U_{t}$. Note that $U$ contains a closed point in $|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ as it is the complement of the union of countably many closed subvarieties and we work over $\mathbb{C}$. Then any $[H]\in U$ satisfies the required conditions. For the general case, let $\overline{X}$ be a projective variety containing $X$ as an open dense subset. Then $\overline{X}\setminus X$ has codimension at least one, and we can assume that there exist irreducible divisors $E_{j}\subset\overline{X}$ for $1\leq j\leq m$ such that $\overline{X}\setminus X\subset\cup_{1\leq j\leq m}E_{j}$. Thus, applying the above result to $\overline{X}$ and $\{E_{j}\}_{1\leq j\leq m}\cup\{D_{i}\}_{i\in\mathbb{N}}$, we get an irreducible curve $\overline{C}\subset\overline{X}$. Then $C:=\overline{C}\cap X$ satisfies all desired properties in the statement. ∎

### v2

@@ 5.3. Families of immersed atomic Lagrangians
Let $p\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to C$ and $q\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to C$ be the relative moduli spaces, which are proper. We denote by $r_{c}\colon M^{\mathcal{Y}_{c}}_{\underline{\sigma^{\prime}}_{c}}(a,b)\to M^{\mathcal{X}_{c}}_{\underline{\sigma}_{c}}(a,b)$ the morphism in Theorem 3.7 for each $c\in C$. By assumption, the push-forward of $\mathcal{Y}\to\mathcal{X}$ and the relative projection functor $\mathrm{D}^{\mathrm{b}}(\mathcal{X})\to\mathcal{K}u(\mathcal{X})$ induce a morphism $r^{\prime}\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ such that $r^{\prime}_{c}=r_{c}$. Therefore, $r^{\prime}$ is proper and quasi-finite, hence finite. Moreover, by [25, Proposition 5.1(1)], the fibers of $p$ are equidimensional of the same dimension and $M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)$ is also equidimensional. By Proposition 4.4, this implies that $r_{0*}\operatorname{\mathcal{O}}_{M^{Y}_{\sigma_{Y}}(a,b)}$ is atomic because $r_{1*}\operatorname{\mathcal{O}}_{M_{\underline{\sigma}_{1}}^{Y^{\prime}}(a,b)}$ is atomic. ∎
@@
5.4. Twisted atomic bundles
In this subsection, we are going to prove Theorem 1.2.
Given a hyper-Kähler manifold $X$ and a Brauer class $\alpha\in\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X}^{*})$, we consider the twisted derived category $\mathrm{D}^{\mathrm{b}}(X,\alpha)$. In [12, Section 4.4], the author extends the theory of atomicity to the case of twisted objects in $\mathrm{D}^{\mathrm{b}}(X,\alpha)$ and proves some properties of twisted atomic sheaves in a specific setting.
On the other hand, for a twisted vector bundle on $X$, the notion of projectively hyperholomorphic bundles is introduced in [75], which has well-behaved deformation properties. We refer to [53, Section 6] for more details.
Proof of Theorem 1.2.
We fix a pair of coprime integers $a,b$ and set $n:=a^{2}+b^{2}+1$. Let $X$ be a general GM fourfold and $Y\subset X$ be a general hyperplane section. By Theorem 5.8, we have an immersed atomic Lagrangian submanifold $r\colon M^{Y}_{\sigma_{Y}}(a,b)\to M^{X}_{\sigma_{X}}(a,b)$. Up to the $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$-action, we may assume that $\sigma_{X}$ and $\sigma_{Y}$ are rational with the same central charge. By Theorem 3.7 and Lemma 3.8, we can choose a non-rigid connected component $L\subset M^{Y}_{\sigma_{Y}}(a,b)$ such that $r$ maps $L$ birationally to its image $Z$. Then, by Lemma 4.9, the restriction of $r$ on $L$, denoted also by $r$, is a connected immersed atomic Lagrangian submanifold.
Recall that $M^{X}_{\sigma_{X}}(a,b)$ forms an open dense subset of the moduli space of projective hyper-Kähler manifolds of degree $22n-$ 2 and divisibility $n-1$ when $X$ varies in an open dense subset of the moduli space of GM fourfolds (cf. [66, Theorem 1.7]). By the Torelli Theorem [32, Theorem 8.1], we can take $X$ such that $\mathrm{H}^{1,1}_{\mathbb{Z}}(M^{X}_{\sigma_{X}}(a,b))=\langle h,f\rangle$, where $h$ is the natural polarization and $f$ induces a Lagrangian fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ such that $q(h,f)=d$ for a positive integer $d$. When $d\gg 0$, the movable cone of $M^{X}_{\sigma_{X}}(a,b)$ equals to the nef cone of $M^{X}_{\sigma_{X}}(a,b)$ by [55, Corollary 2.7] and [61, Lemma 4.3].
Assume that $d$ is sufficiently large and coprime with $n-1=a^{2}+b^{2}$. Then, according to [62, Proposition 5.12] or [52, Theorem 7.13], we can take $X$ such that the fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ is isomorphic to a Tate–Shafarevich twist of $M_{S}(0,\mathcal{L},1-n)\to|\mathcal{L}|$, where $S$ is a K3 surface with $\operatorname{Pic}(S)=\mathbb{Z}\mathcal{L}$. Then, using [12, Theorem 6.2.5], there is a Brauer class $\alpha\in\mathrm{H}^{2}(M^{X}_{\sigma_{X}}(a,b),\operatorname{\mathcal{O}}^{*}_{M^{X}_{\sigma_{X}}(a,b)})$ and a twisted Poincaré sheaf on $M^{X}_{\sigma_{X}}(a,b)\times_{\mathbb{P}^{n}}M_{S}(0,\mathcal{L},1-n)$ such that the corresponding Fourier–Mukai functor $\Phi$ is an equivalence
$\Phi\colon\mathrm{D}^{\mathrm{b}}(M^{X}_{\sigma_{X}}(a,b))\xrightarrow{\simeq}\mathrm{D}^{\mathrm{b}}(M_{S}(0,\mathcal{L},1-n),\alpha).$
In particular, for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$, its image $\Phi(r_{*}\operatorname{\mathcal{F}})\in\mathrm{D}^{\mathrm{b}}(M_{S}(0,\mathcal{L},1-n),\alpha)$ is a twisted atomic object by [12, Proposition 4.4.7].
Now, we claim $f^{n}.r_{*}[L]\neq 0$, which implies that $Z\to\mathbb{P}^{n}$ is finite. Indeed, by Corollary 5.2, we have $h^{\perp}=\ker(r^{*})$. As $q(h,f)=d\neq 0$, we see $f\notin\ker(r^{*})$. By the atomicity, $r^{*}h$ spans $\operatorname{im}(r^{*})$. Then $r^{*}f=t\cdot r^{*}h\in\mathrm{H}^{2}(L,\mathbb{Q})$ for a non-zero $t\in\mathbb{Q}$. Hence, we have $f^{n}.r_{*}[L]=t^{n}\cdot h^{n}.r_{*}[L]\neq 0$.
Since $r$ is finite, $r_{*}\operatorname{\mathcal{F}}$ is a rank one Cohen–Macaulay sheaf on $Z$ for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$ by [40, Corollary 2.73]. Then, from the fact that $Z\to\mathbb{P}^{n}$ is finite, $\Phi(r_{*}\operatorname{\mathcal{F}})$ is a twisted atomic bundle by [12, Proposition 5.2.6, Proposition 6.5.4]. As $Z$ is integral and $r_{*}\operatorname{\mathcal{F}}$ is Cohen–Macaulay of rank one on $Z$, the proof of [12, Theorem 6.5.6] applies in our case. Then we know that $\Phi(r_{*}\operatorname{\mathcal{F}})$ is stable with respect to any suitable polarization.
[…]
Remark 5.12.
We expect that the non-rigid bundle constructed in Theorem 1.2 has a $20$-dimensional deformation space. This can be deduced from the connectedness of $M^{Y}_{\sigma_{Y}}(a,b)$ and the fact that $b_{1}(M^{Y}_{\sigma_{Y}}(a,b))=b_{3}(Y)=20$. We anticipate that the arguments used in [47] for cubic threefolds can be applied in this case as well.
@@ 6.1. Double dual EPW sextics
Recall that by [17, Theorem 3.10], each smooth GM variety $X$ of dimension $\dim X\geq 3$ is uniquely determined by its Lagrangian data
@@
as in [56, Section 2]. See also [17, Appendix B] for an overview. We set $\mathsf{Y}_{A(X)}:=\mathsf{Y}_{A(X)}^{\geq 1}$ and $\mathsf{Y}_{A(X)^{\perp}}:=\mathsf{Y}_{A(X)^{\perp}}^{\geq 1}$, which are called the Eisenbud–Popescu–Walter (EPW) sextic and the dual EPW sextic associated with $X$, respectively.
Moreover, by [58, Section 1.2] and [19, Theorem 5.2], as $A(X)$ does not contain decomposable vectors [^6] This means $\mathbb{P}(A(X))\cap\operatorname{Gr}(3,V_{6})=\varnothing$, which holds for $A(X)$ when $X$ is smooth and $\dim X\geq 3$ by [17, Theorem 3.16]., we have natural double coverings
@@
Since $X$ and $Y$ are both ordinary, by [17, Propoition 3.14], the Lagrangian data of $X$ and $Y$ are $(V_{6},V_{5},A(X))$ and $(V_{6},V_{5},A(Y))$ respectively, with $\dim(A(X)\cap A(Y))=9$. Therefore, part (1) follows directly from the above definition of EPW varieties.
@@
Consequently, $\iota$ is finite and unramified in codimension one. Thus, the natural map $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is also unramified in codimension one. Then, by the smoothness of $\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ and [29, Theorem 1.6], it follows that $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified, and so is $\iota$.
Moreover, since $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified and finite of degree two, by [71, Tag 0B3D] and [71, Tag 00R4], it is flat and hence an étale double covering. In particular, $W$ is smooth. Then, by the uniqueness part of [19, Theorem 5.2], we have $W\cong\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and (2) follows. ∎
@@
It is suggested in [36, Section 5.1] that $i$ is injective, implying that $i$ is an embedding. Nevertheless, this is equivalent to $\varnothing=\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}\cap\mathsf{Y}_{A(X)^{\perp}}^{\geq 2}\subset\mathsf{Y}_{A(X)^{\perp}}$, which is impossible as explained in [24, Lemma 3.2]. See Remark 6.15 for an alternative proof of the non-injectivity of $i$.
@@ 6.1.1. GM threefolds in a fixed GM fourfold
constructed in [49] and [15]. As explained in [36, Section 5.1] (see also [22]), we have a commutative diagram
@@
where the left column is the natural inclusion and the right column is the finite and unramified morphism in Lemma 6.2(2). Using this relation, the following result is obtained in [36, Proposition 5.2].
@@
Since $X$ is very general, by [25, Corollary B.9], there is a commutative diagram
@@
As $X$ and $Y$ are general, we know that $F(X)$ and $F(Y)$ are both smooth by [36, Theorem 3.2] and [49]. Moreover, from [17, Theorem B.2], we may assume that
@@
then $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ are smooth by [19, Theorem 5.2]. Consequently, according to Proposition 6.4, $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an immersed Lagrangian submanifold.
Recall that the moduli stacks of ordinary GM varieties of dimension $n=3$ or $4$ are smooth irreducible separated Deligne–Mumford stacks of finite type over $\mathbb{C}$ (cf. [21, Corollary 5.12] and [41, Proposition A.2]). Since the locus of Hodge-special GM fourfolds is the union of countably many divisors by [16, Corollary 4.6], using Lemma 6.9, we can construct a smooth family of ordinary GM fourfolds $\mathcal{X}\to C$ over a smooth connected curve $C$ and a smooth divisor $\mathcal{Y}\subset\mathcal{X}$ such that the fibers over a point $0\in C$ are $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$. Moreover, there exists an uncountable set $F\subset C$ such that $\mathcal{X}_{c}$ is very general in the sense of Lemma 6.6 for each $c\in F$. Indeed, we start by considering a curve $C^{\prime}$ along with a family $\mathcal{X}^{\prime}\to C^{\prime}$ that satisfies the above assumptions. Next, we lift $C^{\prime}$ to a curve $C$ in the relative Hilbert scheme of hyperplane sections of $\mathcal{X}^{\prime}\to C^{\prime}$, obtaining the family $\mathcal{Y}\subset\mathcal{X}\to C$.
@@
Applying [21, Theorem 3.7, Proposition 4.1] to $\mathcal{X}\to C$ and $\mathcal{Y}\to C$, we obtain two families of Lagrangian data $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{X}))$ and $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{Y}))$ over $C$ [^7] See [21, Definition 3.15] for the precise definition., where $\mathscr{V}_{6}$, $\mathscr{V}_{5}$, $\mathscr{A}(\mathcal{X})$, and $\mathscr{A}(\mathcal{Y})$ are all locally free sheaves on $C$ such that
@@
for each $c\in C$. After replacing $C$ with an open dense neighborhood of $0\in C$, we can assume that these bundles are all trivial. Therefore, as in the proofs of [19, Theorem 5.2] and [18, Proposition 5.27], we get a family $p\colon\widetilde{\mathscr{Y}}_{\mathcal{X}}\to C$ of double dual EPW sextics and a family $q\colon\widetilde{\mathscr{Y}}^{\geq 2}_{\mathcal{Y}}\to C$ of double dual EPW surfaces such that $p^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ and $q^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}$ for each $c\in C$ by replacing $C$ with an open dense neighborhood of $0\in C$. Note that by our assumptions of families $\mathcal{X}$ and $\mathcal{Y}$, both $p$ and $q$ are smooth. We denote by $\mathscr{Y}_{\mathcal{X}}\to C$, $\mathscr{Y}_{\mathcal{Y}}^{\geq 2}\to C$, and $\mathscr{Y}^{\geq 2}_{\mathcal{X}}\to C$ the corresponding families of dual EPW sextics and dual EPW surfaces, respectively. Then $\mathscr{Y}_{\mathcal{X}},\mathscr{Y}^{\geq 2}_{\mathcal{X}}$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}$ are smooth. Note that they are all closed subschemes of $\mathbb{P}(\mathscr{V}^{\vee}_{6})\cong\mathbb{P}(V^{\vee}_{6})\times C$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}_{c}}\subset\mathscr{Y}_{\mathcal{X}_{c}}$ for each $c\in C$, we see $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}\subset\mathscr{Y}_{\mathcal{X}}$. Therefore, by passing to double coverings, taking simultaneous normalization (cf. [39, Theorem 12]), and using a relative version of Lemma 6.2, we have a commutative diagram
@@
Up to now, we have shrunk $C$ near $0\in C$ finitely many times. As $F$ is an infinite set, there is a point $t\in C\cap F$ such that Lemma 6.6 can be applied to $\mathcal{Y}_{t}\hookrightarrow\mathcal{X}_{t}$, i.e. $j_{t}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{t})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{t})^{\perp}}$ is an immersed atomic Lagrangian submanifold. Therefore, using Proposition 4.4, we know that the Lagrangian submanifold $i_{c}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ is atomic for each $c\in C$, and in particular, $j_{0}=i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is atomic. Finally, the computation of the cohomology group is done in [20, Proposition 2.5] and [49, Proposition 0.5], and the orthogonal complement of the kernel of the restriction map is computed in Corollary 5.2 and Remark 3.11. ∎
@@
To this end, we first assume that $X$ is projective. Let $l\colon X\hookrightarrow\mathbb{P}^{n}$ be an embedding, then each hyperplane $H\subset\mathbb{P}^{n}$ intersects with each divisor in $\{D_{i}\}_{i\in\mathbb{N}}$. By [45, Theorem 3.3.1], for each $t\in\mathbb{N}$, there exists an open dense subset $U_{t}\subset|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ such that when $[H]\in U_{t}$, $H\cap X$ and $H\cap D_{i}$ are irreducible and $H$ intersects with $D_{i}$ and $D_{i}\cap D_{j}$ properly for each $i,j\leq t,i\neq j$. Thus, $\dim H\cap D_{i}\cap D_{j}\leq\dim X-3$ for $i,j\leq t,i\neq j$. We define $U:=\cap_{t\in\mathbb{N}}U_{t}$. Note that $U$ contains a closed point in $|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ as it is the complement of the union of countably many closed subvarieties and we work over $\mathbb{C}$. Then any $[H]\in U$ satisfies the required conditions. For the general case, let $\overline{X}$ be a projective variety containing $X$ as an open dense subset. Then $\overline{X}\setminus X$ has codimension at least one, and we can assume that there exist irreducible divisors $E_{j}\subset\overline{X}$ for $1\leq j\leq m$ such that $\overline{X}\setminus X\subset\cup_{1\leq j\leq m}E_{j}$. Thus, applying the above result to $\overline{X}$ and $\{E_{j}\}_{1\leq j\leq m}\cup\{D_{i}\}_{i\in\mathbb{N}}$, we get an irreducible curve $\overline{C}\subset\overline{X}$. Then $C:=\overline{C}\cap X$ satisfies all desired properties in the statement. ∎